Diego Company manufactures one product that is sold for $71 per unit in two geographic regions—the East and West regions. The following information pertains to the company’s first year of operations in which it produced 42,000 units and sold 37,000 units.

Variable costs per unit:
Manufacturing:
Direct materials $ 21
Direct labor $ 12
Variable manufacturing overhead $ 3
Variable selling and administrative $ 5
Fixed costs per year:
Fixed manufacturing overhead $ 840,000
Fixed selling and administrative expense $ 330,000

The company sold 27,000 units in the East region and 10,000 units in the West region. It determined that $160,000 of its fixed selling and administrative expense is traceable to the West region, $110,000 is traceable to the East region, and the remaining $60,000 is a common fixed expense. The company will continue to incur the total amount of its fixed manufacturing overhead costs as long as it continues to produce any amount of its only product.

13. Prepare a contribution format income statement that includes a total column and columns for the East and West regions.
Answer #1
13
Total Company East West
Sales 2627000 1917000 710000
Variable Expenses 1517000 1107000 410000
Contribution Margin 1110000 810000 300000
Traceable Fixed expenses 270000 110000 160000
Region segment margin 840000 700000 140000
Common fixed expenses not traceable to regions 900000
Net operating loss (60000)
Formulas:
East West
Sales =27000*71 =10000*71
Variable Expenses =27000*41 =10000*41
Common fixed expenses not traceable to regions =60000+840000